OpenStudy (anonymous):

What are the similarities and differences between solving the equation, -3x + 2 = 11 and the inequality, -3x + 2 < 11?

OpenStudy (anonymous):

answer please?

OpenStudy (texaschic101):

inequalities are solved just about the same way as equations except for some inequality rules such as when dividing/multiplying by a negative number, it changes the inequality sign.

OpenStudy (anonymous):

that's the answer?

OpenStudy (texaschic101):

equation... -3x + 2 = 11 -3x = 11 - 2 -3x = 9 x = -9/3 x = -3 inequality ... -3x + 2 < 11 -3x < 11 - 2 -3x < 9 x > -9/3 (change the sign when dividing by a negative) x > -3